# Geodesic Synthetic Control Methods for Random Objects and Functional Data

### v1

@@ 6.2. Causal model for geodesic synthetic difference-in-differences
for any $j=1,\dots,J+1$ and any $\bm{\lambda}\in\Delta^{T_{0}-1}$, Assumptions 6.1, 6.2, and 6.3 (i) can be verified.
Figure 7. Illustration of geodesic synthetic difference-in-differences under Assumption 6.2. The circles symbolize objects in the geodesic metric space $\mathcal{M}$. Under Assumption 6.2, $Y_{2:4,1:3}^{(\overline{\bm{w}},\overline{\bm{\lambda}})}=Y_{1,1:3}^{(\overline{\bm{\lambda}})}$ and $Y_{2:4,4:6}^{(\overline{\bm{w}},\oplus)}=Y_{1,4:6}^{(\oplus)}$. The solid red arrow is the geodesic from $Y_{1,1:3}^{(\overline{\bm{\lambda}})}$ to $Y_{1,4:6}^{(\oplus)}(N)$, representing the bias correction for geodesic synthetic controls as effected by geodesic synthetic difference-in-differences.
Figure 8. Illustration of GSDID under Assumption 6.3. The circles symbolize objects in the geodesic metric space $\mathcal{M}$. Arrows depict geodesics connecting them. The dashed red arrow from $Y_{1,1:3}^{(\overline{\bm{\lambda}})}$ to $Y_{1,4:6}^{(\oplus)}(N)$ represents the trend between pre-treatment and post-treatment for unit 1. The solid red arrow represents the parallel trend. The dashed red arrow from $Y_{2:4,4:6}^{(\overline{\bm{w}},\oplus)}$ to $Y_{1,4:6}^{(\oplus)}(N)$ represents the bias correction of the GSC affected by the GSDID.
@@
Theorem 6.1.
Suppose that Assumptions 3.2 and 6.1 hold for the model (6.3). Then under Assumption 6.2 or 6.3, we have
$Y_{1,T_{0}+1:T}^{(\mathrm{GSDID})}=Y_{1,T_{0}+1:T}^{(\oplus)}(N).$ (6.4)
Theorem 6.1 provides a type of double robustness property of the GSDID estimator. Specifically, even if the weights $\overline{\bm{w}}$ for GSC or $\overline{\bm{\lambda}}$ for GDID fail to remove the bias of the method, the combination of weights can compensate for this. This double robustness property corresponds to the double robustness property of synthetic DID for Euclidean outcomes and is similar to the property of the augmented inverse probability weighting estimator for Euclidean outcomes, demonstrating the possibility of trading off the accuracy of the outcome model and the treatment assignment model. We refer to Kurisu et al., (2024) for further discussion of the augmented inverse probability weighting estimator and its double robustness for outcomes taking values in geodesic metric spaces.
Remark 6.2 (Placebo permutation test).
@@
rank of $\frac{\mbox{rank of }d(Y_{1,T_{0}+1:T}^{(\oplus)},Y_{1,T_{0}+1:T}^{(\mathrm{GSDID})})\mbox{ in }\{d(Y_{j,T_{0}+1:T}^{(\oplus)},Y_{j,T_{0}+1:T}^{(\mathrm{GSDID})}):j=1,\ldots,J+1\}}{J+1}.$
Remark 6.3 (Extensions).
@@
$\overline{\bm{\lambda}}_{t}\in\mathop{\rm arg~{}min}\limits_{\bm{\lambda}\in\Delta^{T_{0}-1}}\frac{1}{J}\sum_{j=2}^{J+1}d^{2}(Y_{j,t},Y_{j,1:T_{0}}^{(\bm{\lambda})}),\quad t=T_{0}+1,\dots,T.$
@@
$Y_{1,1:K,1:T_{0}}^{(\bm{\lambda},\oplus)}$ $=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\sum_{t=1}^{T_{0}}\lambda_{t}d^{2}(\nu,Y_{1,1:K,t}^{(\oplus)}),\quad Y_{1,1:K,t}^{(\oplus)}=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\frac{1}{K}\sum_{j=1}^{K}d^{2}(\nu,Y_{1j,t}).$
@@
$Y_{1,1:K,T_{0}+1:T}^{(\oplus,\oplus)}(D)$ $=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\frac{1}{T-T_{0}}\sum_{t=T_{0}+1}^{T}d^{2}(\nu,Y_{1,1:K,t}^{(\oplus)}(D)),\quad D\in\{N,I\},$
$Y_{1,1:K,t}^{(\oplus)}(D)$ $=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\frac{1}{K}\sum_{j=1}^{K}d^{2}(\nu,Y_{1j,t}(N)).$
@@
By appropriately modifying Assumptions 6.16.3, it is also possible to extend Theorem 6.1 to the case with multiple treated units. Further combinations of times and treated units are possible and can be handled analogously.
6.3. Analysis of the effect of the collapse of the Soviet Union on mortality with geodesic synthetic difference-in-differences
The collapse of the Soviet Union in 1991 led to severe economic and social disruptions, profoundly affecting mortality patterns in Russia. The transition from a centrally planned economy to a market-based system was marked by economic instability, rising unemployment, healthcare system deterioration, and increased alcohol consumption, all of which contributed to a dramatic decline in life expectancy. Between 1990 and 1994, life expectancy for Russian men fell from 63.8 to 57.6 years, while for women, it declined from 74.4 to 71.0 years (Notzon et al.,, 1998). Rather than relying solely on summary measures such as life expectancy, a more comprehensive understanding of this demographic shock requires a more detailed analysis, which motivates the analysis of entire age-at-death distributions.
The data supporting such an analysis is available from the Human Mortality Database,, which provides life tables with 5-year time intervals for multiple countries, including Russia and various Western European countries. These life tables are sorted by age and calendar year and can easily be converted into age-at-death distributions. Specifically, for each country, we obtain age-at-death distributions by constructing their probability densities through slightly smoothing the histograms of the age-at-death distribution obtained from the available life table data. For this preprocessing step we use the CreateDensity function from the frechet package (Chen et al., 2023b, ).
We treat Russia as the intervention unit, while the control group consists of 19 Western European countries: Austria, Belgium, Denmark, Finland, France, Germany, Greece, Iceland, Ireland, Italy, Luxembourg, Netherlands, Norway, Portugal, Slovenia, Spain, Sweden, Switzerland, and the United Kingdom. For this analysis, we apply the geodesic synthetic difference-in-differences model (GSDID) as delineated in Sections 6.1 and 6.2. to estimate the causal effect of the Soviet Union’s collapse on Russia’s age-at-death distribution using the Wasserstein metric (see Example 2.4).
Details on the implementation of GSDID are in Algorithm 2. The optimization problems in Steps 1 and 2 of Algorithm 2 are addressed in a similar way to those in Step 1 of Algorithm 1. The third step of Algorithm 2 involves constructing the geodesic transport map, with details provided in Remark 3.5.
@@
$\overline{\bm{w}}\in\mathop{\rm arg~{}min}\limits_{\bm{w}\in\Delta^{J-1}}\frac{1}{T_{0}}\sum_{t=1}^{T_{0}}d^{2}(Y_{1,t},Y_{2:J+1,t}^{(\bm{w})}),$
where $Y_{2:J+1,t}^{(\bm{w})}=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}d^{2}(\nu,Y_{j,t})$ for $t=1,\ldots,T_{0}$;
@@
$\overline{\bm{\lambda}}=\mathop{\rm arg~{}min}\limits_{\bm{w}\in\Delta^{T_{0}-1}}\frac{1}{J}\sum_{j=2}^{J+1}d^{2}(Y_{j,T_{0}+1:T}^{(\oplus)},Y_{j,1:T_{0}}^{(\bm{\lambda})}),$
@@
$Y_{j,T_{0}+1:T}^{(\oplus)}$ $=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\frac{1}{T-T_{0}}\sum_{t=T_{0}+1}^{T}d^{2}(\nu,Y_{j,t}),$
$Y_{j,1:T_{0}}^{(\bm{\lambda})}$ $=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\sum_{t=1}^{T_{0}}\lambda_{t}d^{2}(\nu,Y_{j,t})$
@@
$Y_{2:J+1,1:T_{0}}^{(\overline{\bm{w}},\overline{\bm{\lambda}})}$ $=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}\overline{w}_{j}d^{2}(\nu,Y_{j,1:T_{0}}^{(\overline{\bm{\lambda}})}),$
$Y_{2:J+1,T_{0}+1:T}^{(\overline{\bm{w}},\oplus)}$ $=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}\overline{w}_{j}d^{2}(\nu,Y_{j,T_{0}+1:T}^{(\oplus)});$
@@
We consider four time periods: 1980–1984 and 1985–1989 as pre-treatment periods, and 1990–1994 and 1995–1999 as post-treatment periods, with the Soviet collapse occurring in 1991. The effects of the collapse are expected to be extended over time. Figure 9 illustrates the available data through the estimated probability densities of the age-at-death distributions for males and females in Russia and the control countries. In Russia, a clear leftward shift is observed in the post-treatment periods (1990–1994 and 1995–1999), indicating increased mortality and a decline in longevity. This shift is especially pronounced for males, reflecting a substantial reduction in survival probabilities at younger and middle ages.
@@
Figure 9. Estimated probability densities for age-at-death distributions of females (left) and males (right) for Russia (red) and Western European countries (grey) for the calendar periods 1980–1984, 1985–1989, 1990–1994, and 1995–1999.
@@
When applying GSDID, we reweigh the control countries to ensure that their pre-treatment time trends align with those of Russia, allowing for a difference-in-differences analysis within the geodesic framework. Figure 10 presents the observed and synthetic post-treatment age-at-death densities for both males and females. The GSDID framework captures not only the overall reduction in life expectancy but also the complex age- and gender-specific effects of the economic and societal shocks that resulted from the collapse of the Soviet Union. The results confirm that the collapse caused a significant leftward shift in Russia’s age-at-death distribution, with sharply increased mortality risk across multiple age groups. The effect was particularly severe for males, reflected in a more pronounced leftward shift. A plausible explanation for this gender disparity is the heightened vulnerability of working-age men to economic collapse, labor market instability and increased alcohol-related mortality.
@@
Figure 10. Geodesic synthetic difference-in-differences estimator for age-at-death distributions in Russia, shown separately for females (left) and males (right). The red curves represent the age-at-death density for the synthetic difference-in-differences unit, while the blue curves represent the average post-treatment age-at-death density for Russia.
@@
To assess the statistical significance of this effect, we apply the placebo permutation test described in Remark 6.2. This procedure constructs synthetic age-at-death distributions for each control country by pretending it was treated, allowing us to evaluate whether the observed shift in Russia’s mortality patterns is extreme compared to placebo assignments. The resulting $p$-values for the null hypothesis of no causal effect are 0 for both females and males. These findings provide strong evidence that the collapse of the Soviet Union indeed caused a substantial and statistically significant impact on mortality for nearly all age groups in Russia.
7. Concluding Remarks
There are several directions for future research. Given a specific type of non-Euclidean outcome one often has a choice of several metrics. For example for the case of distributional outcomes, the Wasserstein metric is popular and easy to handle for one-dimensional distributions but it is very tedious to apply it for multivariate distributions (Torous et al.,, 2024) and alternative metrics that are easier to implement for multivariate distributions such as the Fisher-Rao metric or sliced Wasserstein metric (Bonneel et al.,, 2015) may sometimes be preferable. There are also many metrics available for the space of symmetric positive-definite matrices and the space of networks. The specific choice of metric could have a bearing on the results of a causal analysis. Metric selection in general and especially in a causal context remains an open problem (Dubey et al.,, 2024).
Furthermore, it is clearly important to further develop approaches for inference and uncertainty quantification for synthetic control models and their generalizations. One possibility to create an appropriate asymptotic framework is to extend the results of Chernozhukov et al., 2021a ; Chernozhukov et al., 2021b . Another promising avenue for future research is to extend the geodesic synthetic control model to the case of a staggered adoption setting (Ben-Michael et al.,, 2022).
Appendix A Verification of Assumption 3.1
@@ 7. Concluding Remarks
To verify that the causal models for Examples 2.12.5 given in Section 2 satisfy Assumption 3.1, recall that $\overline{\bm{w}}\in\mathop{\rm arg~{}min}\limits_{\bm{w}\in\Delta^{J-1}}\frac{1}{T_{0}}\sum_{t=1}^{T_{0}}d^{2}(Y_{1,t},Y_{2:J+1,t}^{(\bm{w})})$. For $\bm{w}=(w_{2},\dots,w_{J+1})^{\prime}\in\Delta^{J-1}$, define
$U_{2:J+1}^{(\bm{w})}=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}d^{2}(\nu,U_{j}),$
@@
and $\mathbb{W}^{*}=\{\bm{w}\in\Delta^{J-1}:d(U_{1},U_{2:J+1}^{(\bm{w})})=0\}$. Throughout Sections A.1A.5 below, we assume that $\mathbb{W}^{*}$ is not empty.
A.1. Space of networks with the Frobenius metric
@@
$=\mu_{t}+\alpha_{t}\left\{\sum_{j=2}^{J+1}w_{j}U_{j}-\mu_{t}\right\}=\mu_{t}+\alpha_{t}(U_{2:J+1}^{(\bm{w})}-\mu_{t})=g_{t}(U_{2:J+1}^{(\bm{w})}).$ (A.1)
@@
A.2. Space of compositional data with the arc-length distance
@@
Note that $Y_{1,t}=\mathrm{Exp}_{\mu_{t}}(0)=\mu_{t}$. Define $F_{\bm{w}}(\nu)=\sum_{j=2}^{J+1}w_{j}d^{2}(\nu,g_{t}(U_{j})),\bm{w}\in\Delta^{J-1}$ and let $\mathrm{grad}(F_{\bm{w}}(\nu))$ denote the gradient at $\nu\in\mathcal{M}$. Throughout we assume here that among the elements $g_{t}(U_{2}),\dots,g_{t}(U_{J+1})$, at least one is in the interior of $\mathcal{S}_{+}^{d-1}$. Then Theorem 1 in Buss and Fillmore, (2001) implies that Assumption 3.1 (i) holds. For a more general discussion on sufficient conditions for the uniqueness of the Fréchet mean of a distribution on a Riemannian manifold, see Remark 2.8 in Eltzner and Huckemann, (2019) and the references therein. For $\overline{\mathbb{W}}_{t}^{*}=\{\bm{w}\in\Delta^{J-1}:\sum_{j=2}^{J+1}w_{j}\mathrm{Log}_{g_{t}(U_{1})}(g_{t}(U_{j}))=0\}$ one can show that Assumption 3.1 (ii) holds if
$\mathbb{W}_{t}^{*}=\overline{\mathbb{W}}_{t}^{*}.$ (A.3)
@@
=$=\{\bm{w}\in\Delta^{J-1}:A_{t}\sum_{j=2}^{J+1}w_{j}\mathrm{Log}_{U_{1}}(U_{j})=0\}$
=$=\{\bm{w}\in\Delta^{J-1}:\sum_{j=2}^{J+1}w_{j}\mathrm{Log}_{U_{1}}(U_{j})=0\}=\mathbb{W}^{*}.$

### v2

@@ 7 Data Availability Statement
for any $j=1,\dots,J+1$ and any $\bm{\lambda}\in\Delta^{T_{0}-1}$, Assumptions S2, S3, and S4(i) can be verified.
@@
Theorem S1.
Suppose that Assumptions S1 and S2 hold for the model (S3). Then under Assumption S3 or S4, we have
$Y_{1,T_{0}+1:T}^{(\mathrm{GSDID})}=Y_{1,T_{0}+1:T}^{(\oplus)}(N).$ (S4)
Theorem S1 provides a type of double robustness property of the GSDID estimator. Specifically, even if the weights $\overline{\bm{w}}$ for GSC or $\overline{\bm{\lambda}}$ for GDID fail to remove the bias of the method, the combination of weights can compensate for this. This double robustness property corresponds to the double robustness property of synthetic DID for Euclidean outcomes and is similar to the property of the augmented inverse probability weighting estimator for Euclidean outcomes, demonstrating the possibility of trading off the accuracy of the outcome model and the treatment assignment model. We refer to Kurisu et al. (2024) for further discussion of the augmented inverse probability weighting estimator and its double robustness for outcomes taking values in geodesic metric spaces.
Remark S3 (Placebo permutation test).
@@
$\frac{\mbox{rank of }d(Y_{1,T_{0}+1:T}^{(\oplus)},Y_{1,T_{0}+1:T}^{(\mathrm{GSDID})})\mbox{ in }\{d(Y_{j,T_{0}+1:T}^{(\oplus)},Y_{j,T_{0}+1:T}^{(\mathrm{GSDID})}):j=1,\ldots,J+1\}}{J+1}.$
Remark S4 (Extensions).
@@
$\overline{\bm{\lambda}}_{t}\in\operatorname*{arg\,min}_{\bm{\lambda}\in\Delta^{T_{0}-1}}\frac{1}{J}\sum_{j=2}^{J+1}d^{2}(Y_{j,t},Y_{j,1:T_{0}}^{(\bm{\lambda})}),\quad t=T_{0}+1,\dots,T.$
@@
$Y_{1,1:K,1:T_{0}}^{(\bm{\lambda},\oplus)}$ $=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{t=1}^{T_{0}}\lambda_{t}d^{2}(\nu,Y_{1,1:K,t}^{(\oplus)}),\quad Y_{1,1:K,t}^{(\oplus)}=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\frac{1}{K}\sum_{j=1}^{K}d^{2}(\nu,Y_{1j,t}).$
@@
$Y_{1,1:K,T_{0}+1:T}^{(\oplus,\oplus)}(D)$ $=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\frac{1}{T-T_{0}}\sum_{t=T_{0}+1}^{T}d^{2}(\nu,Y_{1,1:K,t}^{(\oplus)}(D)),\quad D\in\{N,I\},$
$Y_{1,1:K,t}^{(\oplus)}(D)$ $=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\frac{1}{K}\sum_{j=1}^{K}d^{2}(\nu,Y_{1j,t}(N)).$
@@
By appropriately modifying Assumptions S2S4, it is also possible to extend Theorem S1 to the case with multiple treated units. Further combinations of times and treated units are possible and can be handled analogously.
S.2.3 Analysis of the effect of the collapse of the Soviet Union on mortality with geodesic synthetic difference-in-differences
The collapse of the Soviet Union in 1991 led to severe economic and social disruptions, profoundly affecting mortality patterns in Russia. The transition from a centrally planned economy to a market-based system was marked by economic instability, rising unemployment, healthcare system deterioration, and increased alcohol consumption, all of which contributed to a dramatic decline in life expectancy. Between 1990 and 1994, life expectancy for Russian men fell from 63.8 to 57.6 years, while for women, it declined from 74.4 to 71.0 years (Notzon et al., 1998). Rather than relying solely on summary measures such as life expectancy, a more comprehensive understanding of this demographic shock requires a more detailed analysis, which motivates the analysis of entire age-at-death distributions.
The data supporting such an analysis is available from the Human Mortality Database, which provides life tables with 5-year time intervals for multiple countries, including Russia and various Western European countries. These life tables are sorted by age and calendar year and can easily be converted into age-at-death distributions. Specifically, for each country, we obtain age-at-death distributions by constructing their probability densities through slightly smoothing the histograms of the age-at-death distribution obtained from the available life table data. For this preprocessing step we use the CreateDensity function from the frechet package (Chen et al., 2023b).
We treat Russia as the intervention unit, while the control group consists of 19 Western European countries: Austria, Belgium, Denmark, Finland, France, Germany, Greece, Iceland, Ireland, Italy, Luxembourg, Netherlands, Norway, Portugal, Slovenia, Spain, Sweden, Switzerland, and the United Kingdom. For this analysis, we apply the geodesic synthetic difference-in-differences model (GSDID) as delineated in Sections S.2.1 and S.2.2. to estimate the causal effect of the Soviet Union’s collapse on Russia’s age-at-death distribution using the Wasserstein metric (see Example 4).
@@
$\overline{\bm{w}}\in\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J-1}}\frac{1}{T_{0}}\sum_{t=1}^{T_{0}}d^{2}(Y_{1,t},Y_{2:J+1,t}^{(\bm{w})}),$
where $Y_{2:J+1,t}^{(\bm{w})}=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}d^{2}(\nu,Y_{j,t})$ for $t=1,\ldots,T_{0}$;
@@
$\overline{\bm{\lambda}}=\operatorname*{arg\,min}_{\bm{\lambda}\in\Delta^{T_{0}-1}}\frac{1}{J}\sum_{j=2}^{J+1}d^{2}(Y_{j,T_{0}+1:T}^{(\oplus)},Y_{j,1:T_{0}}^{(\bm{\lambda})}),$
@@
$Y_{j,T_{0}+1:T}^{(\oplus)}$ $=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\frac{1}{T-T_{0}}\sum_{t=T_{0}+1}^{T}d^{2}(\nu,Y_{j,t}),$
$Y_{j,1:T_{0}}^{(\bm{\lambda})}$ $=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{t=1}^{T_{0}}\lambda_{t}d^{2}(\nu,Y_{j,t})$
@@
$Y_{2:J+1,1:T_{0}}^{(\overline{\bm{w}},\overline{\bm{\lambda}})}$ $=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}\overline{w}_{j}d^{2}(\nu,Y_{j,1:T_{0}}^{(\overline{\bm{\lambda}})}),$
$Y_{2:J+1,T_{0}+1:T}^{(\overline{\bm{w}},\oplus)}$ $=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}\overline{w}_{j}d^{2}(\nu,Y_{j,T_{0}+1:T}^{(\oplus)});$
@@
Details on the implementation of GSDID are in Algorithm 2. The optimization problems in Steps 1 and 2 of Algorithm 2 are addressed in a similar way to those in Step 1 of Algorithm 1. The third step of Algorithm 2 involves constructing the geodesic transport map, with details provided in Section S.1.1.
@@
Figure S4: Estimated probability densities for age-at-death distributions of females (left) and males (right) for Russia (red) and Western European countries (grey) for the calendar periods 1980–1984, 1985–1989, 1990–1994, and 1995–1999.
We consider four time periods: 1980–1984 and 1985–1989 as pre-treatment periods, and 1990–1994 and 1995–1999 as post-treatment periods, with the Soviet collapse occurring in 1991. The effects of the collapse are expected to be extended over time. Figure S4 illustrates the available data through the estimated probability densities of the age-at-death distributions for males and females in Russia and the control countries. In Russia, a clear leftward shift is observed in the post-treatment periods (1990–1994 and 1995–1999), indicating increased mortality and a decline in longevity. This shift is especially pronounced for males, reflecting a substantial reduction in survival probabilities at younger and middle ages.
@@
When applying GSDID, we reweigh the control countries to ensure that their pre-treatment time trends align with those of Russia, allowing for a difference-in-differences analysis within the geodesic framework. Figure S5 presents the observed and synthetic post-treatment age-at-death densities for both males and females. The GSDID framework captures not only the overall reduction in life expectancy but also the complex age- and gender-specific effects of the economic and societal shocks that resulted from the collapse of the Soviet Union. The results confirm that the collapse caused a significant leftward shift in Russia’s age-at-death distribution, with sharply increased mortality risk across multiple age groups. The effect was particularly severe for males, reflected in a more pronounced leftward shift. A plausible explanation for this gender disparity is the heightened vulnerability of working-age men to economic collapse, labor market instability and increased alcohol-related mortality.
@@
Figure S5: Geodesic synthetic difference-in-differences estimator for age-at-death distributions in Russia, shown separately for females (left) and males (right). The red curves represent the age-at-death density for the synthetic difference-in-differences unit, while the blue curves represent the average post-treatment age-at-death density for Russia.
@@
To assess the statistical significance of this effect, we apply the placebo permutation test described in Remark S3. This procedure constructs synthetic age-at-death distributions for each control country by pretending it was treated, allowing us to evaluate whether the observed shift in Russia’s mortality patterns is extreme compared to placebo assignments. The resulting $p$-values for the null hypothesis of no causal effect are 0 for both females and males. These findings provide strong evidence that the collapse of the Soviet Union indeed caused a substantial and statistically significant impact on mortality for nearly all age groups in Russia.
S.3 Optimization Details
This section provides additional computational details for the implementation of geodesic synthetic control (GSC) and geodesic synthetic difference-in-differences (GSDID) estimators. We present explicit formulations of the optimization problems, describe how weighted Fréchet means are computed in different geodesic metric spaces, and discuss the solver strategies adopted in each case.
S.3.1 Unit and time weights
Both GSC and GSDID rely on optimizing weights attached to observational indices rather than features of the objects themselves. Specifically, the unit weights $\overline{\bm{w}}\in\Delta^{J-1}$ form convex combinations of control units, while the time weights $\overline{\bm{\lambda}}\in\Delta^{T_{0}-1}$ form convex combinations of pre-treatment periods. Random objects are then averaged via weighted Fréchet means in the respective metric space. Thus, the weights govern the influence of units or time periods but do not vary across internal coordinates of the objects (such as quantiles of a distribution). This mirrors the SDID construction and preserves interpretability in the geodesic extension.
S.3.2 General optimization problems
The estimation of unit and time weights in GSC and GSDID is formulated as minimization problems involving distances between the treated unit and weighted Fréchet means of control units.
For the GSC estimator, only unit weights are involved. Specifically, the unit weights $\overline{\bm{w}}\in\Delta^{J-1}$ are chosen to minimize the average squared distance between the treated outcome $Y_{1,t}$ and the synthetic control outcome $Y^{(\bm{w})}_{2:J+1,t}$ over the pre-treatment period,
$\overline{\bm{w}}\in\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J-1}}\frac{1}{T_{0}}\sum_{t=1}^{T_{0}}d^{2}\!\Bigl(Y_{1,t},\,Y^{(\bm{w})}_{2:J+1,t}\Bigr)$,
where the synthetic control at time $t$ is defined through the weighted Fréchet mean
$Y^{(\bm{w})}_{2:J+1,t}=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}d^{2}(\nu,Y_{j,t}).$
For the GSDID estimator, both unit and time weights are estimated. The unit weights $w$ are obtained in the same way as for GSC, while the time weights $\overline{\bm{\lambda}}\in\Delta^{T_{0}-1}$ are determined by matching the pre-treatment and post-treatment trajectories of the control units,
$\overline{\bm{\lambda}}\in\operatorname*{arg\,min}_{\bm{\lambda}\in\Delta^{T_{0}-1}}\frac{1}{J}\sum_{j=2}^{J+1}d^{2}\!\Bigl(Y^{(\oplus)}_{j,T_{0}+1:T},\,Y^{(\bm{\lambda})}_{j,1:T_{0}}\Bigr)$,
where $Y^{(\oplus)}_{j,T_{0}+1:T}$ denotes the post-treatment Fréchet mean of unit $j$ and $Y^{(\bm{\lambda})}_{j,1:T_{0}}$ the weighted pre-treatment Fréchet mean.
In both cases, the optimization is carried out over the probability simplex, with the geometry of the underlying metric space determining how the Fréchet means are evaluated.
S.3.3 Spaces with closed-form weighted Fréchet means
For networks (Example 1), symmetric positive-definite matrices (Example 3), distributions (Example 4), and functional data (Example 5), the weighted Fréchet mean admits an explicit expression under the metrics considered in this paper. In these cases, the outer optimization problems in Section S.3.2 reduce to convex quadratic programs in $w$ (and $\lambda$), which can be solved efficiently using standard quadratic programming solvers such as OSQP (Stellato et al., 2020). This yields numerically stable and fast implementations.
S.3.4 Compositional data
For compositional data (Example 2), the situation is more complex. When mapped to the positive orthant of the sphere, the weighted Fréchet mean has no closed form and must be obtained via iterative algorithms on the manifold. Each evaluation of the objective therefore requires solving an inner Fréchet mean problem, which creates a bilevel optimization problem. Because analytical differentiation is intractable in this nonlinear setting, we employ a derivative-free constrained optimization strategy, specifically the COBYLA algorithm (Powell, 1994) as implemented in the nloptr package (Johnson, 2008). The inner Fréchet mean computations are carried out using the manifold package (Dai et al., 2021).
S.3.5 Practical remarks
For the tractable cases with closed-form weighted Fréchet means, optimization is straightforward and stable. For compositional data, convergence of the iterative mean calculation can be slow when many observations lie close to zero components, in which case regularization (e.g., adding a small constant to proportions before normalization) improves stability.
S.4 Verification of Assumption 3.1
@@
To verify that the causal models for Examples 1–5 given in Section 2 satisfy Assumption 3.1, recall that $\overline{\bm{w}}\in\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J-1}}\frac{1}{T_{0}}\sum_{t=1}^{T_{0}}d^{2}(Y_{1,t},Y_{2:J+1,t}^{(\bm{w})})$. For $\bm{w}=(w_{2},\dots,w_{J+1})^{\prime}\in\Delta^{J-1}$, define
$U_{2:J+1}^{(\bm{w})}=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}d^{2}(\nu,U_{j}),$
@@
and $\mathbb{W}^{*}=\{\bm{w}\in\Delta^{J-1}:d(U_{1},U_{2:J+1}^{(\bm{w})})=0\}$. Throughout Sections S.4.1S.4.5 below, we assume that $\mathbb{W}^{*}$ is not empty.
S.4.1 Space of networks with the Frobenius metric
@@
$=\mu_{t}+\alpha_{t}\left\{\sum_{j=2}^{J+1}w_{j}U_{j}-\mu_{t}\right\}=\mu_{t}+\alpha_{t}(U_{2:J+1}^{(\bm{w})}-\mu_{t})=g_{t}(U_{2:J+1}^{(\bm{w})}).$ (S5)
@@
S.4.2 Space of compositional data with the arc-length distance
@@
Note that $Y_{1,t}=\mathrm{Exp}_{\mu_{t}}(0)=\mu_{t}$. Define $F_{\bm{w}}(\nu)=\sum_{j=2}^{J+1}w_{j}d^{2}(\nu,g_{t}(U_{j})),\bm{w}\in\Delta^{J-1}$ and let $\mathrm{grad}(F_{\bm{w}}(\nu))$ denote the gradient at $\nu\in\mathcal{M}$. Throughout we assume here that among the elements $g_{t}(U_{2}),\dots,g_{t}(U_{J+1})$, at least one is in the interior of $\mathcal{S}_{+}^{d-1}$. Then Theorem 1 in Buss and Fillmore (2001) implies that Assumption 3.1(i) holds. For a more general discussion on sufficient conditions for the uniqueness of the Fréchet mean of a distribution on a Riemannian manifold, see Remark 2.8 in Eltzner and Huckemann (2019) and the references therein. For $\overline{\mathbb{W}}_{t}^{*}=\{\bm{w}\in\Delta^{J-1}:\sum_{j=2}^{J+1}w_{j}\mathrm{Log}_{g_{t}(U_{1})}(g_{t}(U_{j}))=0\}$ one can show that Assumption 3.1(ii) holds if
$\mathbb{W}_{t}^{*}=\overline{\mathbb{W}}_{t}^{*}.$ (S7)
@@
$=\{\bm{w}\in\Delta^{J-1}:A_{t}\sum_{j=2}^{J+1}w_{j}\mathrm{Log}_{U_{1}}(U_{j})=0\}$
$=\{\bm{w}\in\Delta^{J-1}:\sum_{j=2}^{J+1}w_{j}\mathrm{Log}_{U_{1}}(U_{j})=0\}=\mathbb{W}^{*}.$